Ordered: IV of 1000ml to infuse at 125ml/hr. The drop factor is 20. How many gtt/min will you set IV to infuse?
Answer #1

ANSWER.

* Calculation of gtt/min.

           gtt/min = Volume(ml) / Time(min) * Drop factor(gtt/ml).

Here

       Volume(ml) = 125 ml

       Time(min) = 60 min ( 1hr).

       Drop factor = 20 gtts/ml.

As per above formula

          Volume(ml) of drug required = D/H * Q

                                                        = 125/60 * 20

                                                        = 2.08 * 20

                                                        = 41.6 gtt/min.